# Enhancing (quasi-)long-range order in a two-dimensional driven crystal

### v1

@@ Abstract
It has been recently shown that 2D systems can exhibit crystalline phases with long-range translational order showcasing a striking violation of the Hohenberg-Mermin-Wagner (HMW) theorem which is valid at equilibrium. This is made possible by athermal driving mechanisms that inject energy into the system without exciting long wavelength modes of the density field. However, as thermal fluctuations are superimposed on the non-equilibrium driving, long-range translational order is inevitably lost. In this paper, we discuss the possibility of exploiting non-equilibrium effects to suppress arbitrarily large density fluctuations even when a global thermal bath is coupled to the system. We introduce a model of a harmonic crystal driven both by a global thermal bath and by a momentum conserving noise, where the typical observables related to density fluctuations and long-range translational order can be analytically derived and put in relation. This model allows us to rationalize the violation of the HMW theorem observed in previous studies through the prediction of large-wavelength phonons which thermalize at a vanishing effective temperature when the global bath is switched off. The conceptual framework introduced through this theory is then applied to numerical simulations of a hard-disk solid in contact with a thermal bath and driven out-of-equilibrium by active collisions. Our numerical analysis demonstrates how varying driving and dissipative parameters can lead to an arbitrary enhancement of the quasi-long-range order in the system regardless of the applied global noise amplitude. Finally, we outline a possible experimental procedure to apply our results to a realistic granular system.
@@ I Introduction
While the HMW theorem arises from equilibrium statistical physics, the majority of systems encountered in nature operate out-of-equilibrium. Consequently, the applicability of the theorem to such systems is not assured. As an example, a notable breakdown is observed in the flocking behavior of birds [9]. This process can be modeled by an active $XY$ model – the Vicsek model [10], showcasing spin-spin long-range order in 2D [11, 12, 13, 14, 15], in clear violation of the HMW theorem. Various studies have further explored potential violations of this theorem in non-equilibrium spin-like systems due to forcing at multiple temperatures [16, 17]; advection of the order parameter through a shear flow [18, 19, 20, 21, 22]; or a colored noise [23, 24]. These ingredients have been shown to be crucial in order to obtain spin-spin long-range order. However, until recently, the prerequisites for translational long-range order in 2D crystals remained unknown. Indeed, while exceptions were found in nematic crystals [25] or chiral active matter [26], Mermin-Wagner fluctuations are usually found to be enhanced in the presence of active forces [27, 28, 29] and as a result, an increase of the lower critical dimension is often observed due to giant number fluctuations [30, 31].
However, Galliano et al. presented compelling evidence of the breakdown of the HMW theorem in 2D crystalline systems [32] formed in the active state of a random organization model. The authors showed that, unlike in short-ranged equilibrium systems, a simple non-equilibrium hyperuniform crystal exhibits translational long-range order, as already hinted in Refs. 33, 34, due to the absence of thermal fluctuations at large length scale.
@@ II.1 The model
The first term on the right-hand side of the equation represents the harmonic interaction between neighbors with spring constant $K$.
@@
Finally, the last term in Eq. (1) represents a second bath, at temperature $T_{loc}$, conserving the momentum and the position of the COM. To fulfill this requirement, we use a discretized version of Model B [47, 48, 49, 50, 51, 52]:
$\bm{F}_{loc}=-\gamma_{loc}\sum_{\{\bar{n}\}}({\bm{\dot{u}}}_{n}-\bm{\dot{u}}_{\bar{n}})+\sqrt{2\gamma_{loc}T_{loc}}(\bm{\nabla}\cdot\bm{\Xi}_{{n}}).$ (4)
The conservation of $\sum_{n}\dot{\bm{u}}_{n}$ is ensured by a discrete Laplacian of $\dot{\bm{u}}$ for the damping $\gamma_{loc}$, taken to be the average over first neighbors and a divergence acting on a rank 2 random tensor $\bm{\Xi}$. This type of damping naturally arises in granular gases where collisions tend to align particles [53], in active matter with effective alignment [54, 46] or in Dissipative Particle Dynamics [55, 56, 57]. More generally, with the corresponding equilibrium noise, it plays the role of a discrete hydrodynamic viscosity. Note that if every particle moves in the same direction, neither damping nor noise is applied, in line with the idea that this local noise might arise from collisions between particles and acts locally. Our model constitutes a simplified lattice equivalent to a fluctuating hydrodynamics description of a solid [58, 59, 60, 61, 62, 63, 64]. The variance and the average of the rank 2 tensor $\bm{\Xi}_{n}$ are fully determined by our assumption that the FDT holds separately for both noises [42, 65]:
$\langle\bm{\Xi}_{n}^{\alpha,\mu}(t)\bm{\Xi}_{m}^{\beta,\nu}(t^{\prime})\rangle=\delta(t^{\prime}-t)\delta^{\alpha,\beta}\delta^{\mu,\nu}\delta_{n,m},$ (5)
We further discuss the peculiarity of the noise as well as the expression of the discrete Laplacian and divergence in Appendix A following Ref. 66.
@@
Since the dispersion relation must vanishes at low $k$: $\omega_{k}^{2}\sim(a\bm{k})^{2}$ [68] we obtain the two limiting cases:
@@
That is, the large length scale temperature is controlled by $T_{com}$ since the energy created on small length scales by the local bath is damped by $\gamma_{com}$ on every scale. While the small length scale temperature is controlled by both the global and local bath. In the following, we will assume that the system is more weakly coupled to the local bath compared to the global bath: $\gamma_{loc}\gg\gamma_{com}$ which leads to $T_{k\to\infty}=T_{loc}$.
The equilibrium-like nature of our equation stems from the assumption that both baths are delta correlated and that each mode is independent on the others in $k$ space. This would not be the case in a realistic system with anharmonic terms for instance. Nonetheless, energy injection would still be localized at the scale of collisions – albeit with an energy flux between modes [69] – and large wavelength thermal fluctuations would remain damped by $\gamma_{com}$. Therefore, the rough phenomenology introduced in this section is expected to hold as well, in these more complex cases.
@@ II.2 Hyperuniformity
As already noted in Ref. 32, a key requirement to obtain a long-range ordered 2D crystal is hyperuniformity i.e. the suppression of density fluctuations at arbitrarily large length scales. Within the context of the proposed theoretical model, we derive here the analytical expression of the structure factor $S(\bm{k})$ whose vanishing low-$\bm{k}$ limit indicates the presence of hyperuniformity in a system [33].
@@
When $T_{com}=0$ and $\gamma_{com}\neq 0$, since the SDF is finite at low $k$, the structure factor behaves as $S(\bm{k})\sim\bm{k}^{2}$, unveiling the hyperuniformity of our non-equilibrium system when the COM position is conserved by the noise at finite global damping. This aligns with results found in Refs. 26, 50, 72, 73, 74, 32, 75 and confirms that our model includes a suitable limit to study the emergence of long-range order in dimension lower than three.
@@
where $\bm{G}$ is one of the inner Bragg-peak vectors of the crystal and the last equality follows from the gaussianity of the stationary probability distribution function. In a long-range ordered crystal, the correlation function decays to a constant value while in a quasi-long-range 2D equilibrium crystal it exhibits an algebraic decay [4, 7, 77, 78, 79, 80]. From Eq. (14), it is clear that the asymptotic behavior of the displacement correlation function $\langle(\bm{u}_{n}-\bm{u}_{m})^{2}\rangle$ governs the translational long-range behavior of the system. Specifically, if the displacement correlation function reaches a constant or decays at large distances, true long-range order is expected, otherwise, only quasi-long-range order or short-range order is found. In the infinite system size limit and at large distances compared to the lattice spacing it can be approximated as (see Appendix C) :
@@
where the angle average of the dot product appearing in Eq. (14) was taken into account by assuming an isotropic crystal. At non zero $T_{com}$, the large-scale decay of the correlation function is algebraic indicating quasi-long-range order [81]. Nonetheless, in contrast to the equilibrium picture, the average kinetic energy of the system can be arbitrarily high, the only crucial factor for the translational quasi-long-range behavior of the system at large scale, is the temperature of the COM or equivalently, the temperature of the phonons existing on these scales. We also note that, as expected when the COM is conserved at finite global damping ($T_{com}=0$ and $\gamma_{com}\neq 0$), the system displays genuine long-range order as observed in Ref. 32 and the HMW theorem is broken.
@@
with $\beta>0$.
@@ II.4 Breakdown of the HMW theorem
It is interesting to note that hyperuniformity of equilibrium crystals is solely achieved through long-range or external interactions [33], in contrast with our system and that of others, where hyperuniformity is allowed through the breakdown of the FDT and correlated noise [24, 82, 26, 83, 84]. This is reminiscent of long-range correlations induced by bulk conservation law and dissipation in self-organized criticality [85, 86, 44, 53, 87, 88, 89, 90, 91, 92]. See Ref. 93 for a recent review. Moreover, COM conservation is not the only route to hyperuniformity and systems with multiplicative [94, 82] or time correlated noise can have suppressed long-range fluctuations [35, 24]. Additionally, recent work on particular non-equilibrium collective excitations in active solids, known as entropons [67, 95], highlight that on a general basis, active forces only add an additional term to the SDF, linked to entropy production but do not affect its phononic part. Thus, active forces alone cannot tame the $1/k^{2}$ divergence of the phononic part of the equilibrium SDF and in general, increase the density fluctuations [29, 27, 28]. Indeed, active short-ranged systems exhibiting hyperuniformity usually lose this property in the presence of a global thermal equilibrium-like noise [96, 37, 97, 36]. This supports the idea that non-equilibrium-like colored or multiplicative noises are crucial for hyperuniformity and translational long-range order in 2D crystals featuring short-ranged interactions. Further discussions about entropons in our system can be found in Appendix D. Note however, that hyperuniformity in passive or active scalar field theories has been found even in the presence of a thermal noise for model B type equations [98, 99, 100, 101].
@@ III.1 A coarse grained model
We explore a practical application of our theoretical framework to investigate the breakdown of the HMW theorem and the dependence of (quasi-)long-range order on the noise in realistic non-equilibrium systems. To conduct this investigation, we use the model introduced in Ref. 38 with an additional global noise. This can be thought as an underdamped continuous-time analogue of a random organization model [102] describing sheared suspensions [103]. In the absence of global noise, these models present an absorbing phase transition, are hyperuniform [33] at the transition point [73, 82] and, if the COM is conserved, in the active state [72, 75, 26]. As will be further discussed in the conclusion, this model has been used to describe the experimental realization of vibrated quasi-2d granular systems [104, 38, 105].
We study this system by performing hybrid event-driven/time-stepped molecular dynamics simulations [106] (see Appendix E) of $N$ active hard disks of diameter $\sigma$ and mass $m$ in a 2D square of size $L$ with periodic boundary conditions. We define an arbitrary unit of time $\hat{\tau}$. The (granular) temperature $T$ of the system is defined as its global kinetic energy:
@@
where $0\leq\alpha\leq 1$ is the coefficient of restitution, $\Delta>0$ is a velocity injection term, $\bm{v}_{i}^{\prime}$ the post-collision velocity of particle $i$, $\bm{v}_{i}$ its pre-collision velocity, and $\hat{\bm{\sigma}}_{ij}$ and $\bm{v}_{ij}$ are respectively the unit vector joining particles $i$ and $j$ and the relative velocity between them. In the limit $\Delta=0$, the usual collision rule of dissipative hard disks is recovered.
@@
Hence, to recapitulate the differences and similarities between the harmonic crystal presented in the theory and the granular hard disks of interest here: both models feel an external global Gaussian bath at temperature $T_{com}$ and damping $\gamma_{com}$, however, the active collisions in the granular system are modeled as an effective local bath at temperature $T_{loc}$ and damping $\gamma_{com}\omega_{k}^{2}$ in the theory. From a physical standpoint, the effective $T_{loc}$ in the granular system is technically a function of every parameter of our simulation including, $T_{com}$ and $\gamma_{com}$ because the energy change at the collision depends on the relative velocities of the particles colliding, as understood from Eq. (23), which are affected by the global bath and the packing fraction. This will not change the physical picture given in this article nonetheless, we should keep it in mind. This peculiarity is further discussed in Appendix F.
When $T_{com}=0$, the dynamics given by Eq. (23) and (21) conserves the position of the COM if the initial total momentum of the system is equal to 0 (otherwise, it exponentially relaxes to 0). Hyperuniformity is thus expected [26] and we recover the system studied in Ref. 38 with an absorbing state at low density (a feature absent in the lattice model because $T_{loc}$ is independent on $T_{com}$ and $\gamma_{com}$). Indeed, at low density or packing fraction $\phi=N\pi\sigma^{2}/4L^{2}$, the active collisions cannot compensate for the energy dissipated by the drag during the free flight leading the system to an arrested state. Denser systems attain an active steady state where the energy injection at collision is compensated by the dissipation of the damping.

### v2

@@ Abstract
It has been recently shown that 2D systems can exhibit crystalline phases with long-range translational order showcasing a striking violation of the Hohenberg-Mermin-Wagner (HMW) theorem which is valid at equilibrium. This is made possible by athermal driving mechanisms that inject energy into the system without exciting long wavelength modes of the density field, thereby inducing hyperuniformity. However, as thermal fluctuations are superimposed on the non-equilibrium driving, long-range translational order is inevitably lost. Here, we discuss the possibility of exploiting non-equilibrium effects to suppress arbitrarily large density fluctuations even when a global thermal bath is coupled to the system. We introduce a model of a harmonic crystal driven both by a global thermal bath and by a momentum conserving noise, where the typical observables related to density fluctuations and long-range translational order can be analytically derived and put in relation. This model allows us to rationalize the violation of the HMW theorem observed in previous studies through the prediction of large-wavelength phonons which thermalize at a vanishing effective temperature when the global bath is switched off. The conceptual framework introduced through this theory is then applied to numerical simulations of a hard-disk solid in contact with a thermal bath and driven out-of-equilibrium by active collisions. Our numerical analysis demonstrates how varying driving and dissipative parameters can lead to an arbitrary enhancement of the quasi-long-range order in the system regardless of the applied global noise amplitude. Finally, we outline a possible experimental procedure to apply our results to a realistic granular system.
@@ I Introduction
While the HMW theorem arises from equilibrium statistical physics, the majority of systems encountered in nature operate out-of-equilibrium. Consequently, the applicability of the theorem to such systems is not assured. As an example, a notable breakdown is observed in the flocking behavior of birds [9]. This process can be modeled by an active $XY$ model – the Vicsek model [10], showcasing spin-spin long-range order in 2D [11, 12, 13, 14, 15], in clear violation of the HMW theorem. Various studies have further explored potential violations of this theorem in non-equilibrium spin-like systems due to forcing at multiple temperatures [16, 17]; advection of the order parameter through a shear flow [18, 19, 20, 21, 22]; or a colored noise [23, 24]. These ingredients have been shown to be crucial in order to obtain spin-spin long-range order which, in turn, can be easily destabilized by other out-of-equilibrium effects [25]. However, until recently, the prerequisites for translational long-range order in 2D crystals remained unknown. Indeed, while exceptions were found in nematic crystals [26] or chiral active matter [27], Mermin-Wagner fluctuations are usually found to be enhanced in the presence of active forces [28, 29, 30] and as a result, an increase of the lower critical dimension is often observed due to giant number fluctuations [31, 32].
However, Galliano et al. presented compelling evidence of the breakdown of the HMW theorem in 2D crystalline systems [33] formed in the active state of a random organization model. The authors showed that, unlike in short-ranged equilibrium systems, a simple non-equilibrium hyperuniform crystal exhibits translational long-range order, as already hinted in Refs. 34, 35, due to the absence of thermal fluctuations at large length scale.
@@ II.1 The model
The first term on the right-hand side of the equation represents the harmonic interaction between nearest neighbors.
@@
Finally, the last term in Eq. (1) represents a second bath, at temperature $T_{loc}$, conserving the momentum and the position of the COM. To fulfill this requirement, we use a discretized version of Model B [47, 48, 49, 50, 51, 52, 53]:
$\bm{F}_{loc}=-\gamma_{loc}\sum_{\{\bar{n}\}}({{\dot{\bm{u}}}}_{n}-{\dot{\bm{u}}}_{\bar{n}})+\sqrt{2\gamma_{loc}T_{loc}}(\bm{\nabla}\cdot\bm{\Xi}_{{n}}).$ (4)
The conservation of $\sum_{n}\dot{\bm{u}}_{n}$ is ensured by a discrete Laplacian of $\dot{\bm{u}}$ for the damping $\gamma_{loc}$, taken to be the average over first neighbors and a discrete divergence acting on a rank 2 random tensor $\bm{\Xi}$. This type of damping naturally arises in granular gases where collisions tend to align particles [54], in active matter with effective alignment [55, 46] or in Dissipative Particle Dynamics [56, 57, 58]. More generally, with the corresponding equilibrium noise, it plays the role of a discrete hydrodynamic viscosity and can be derived from the Mori-Zwanzig formalism [59]. Note that if every particle moves in the same direction, neither damping nor noise is applied, in line with the idea that this local noise might arise from collisions between particles and acts locally. Our model constitutes a simplified lattice equivalent to a fluctuating hydrodynamics description of a solid [60, 61, 62, 63, 64, 65, 66]. The variance and the average of the rank 2 tensor $\bm{\Xi}_{n}$ are fully determined by our assumption that the FDT holds separately for both noises [42, 67]:
$\langle\Xi_{n}^{\alpha,\mu}(t)\Xi_{m}^{\beta,\nu}(t^{\prime})\rangle=\delta(t^{\prime}-t)\delta^{\alpha,\beta}\delta^{\mu,\nu}\delta_{n,m},$ (5)
We further discuss the peculiarity of the noise as well as the expression of the discrete Laplacian and divergence in Appendix A following Ref. 68.
@@
Since the dispersion relation must vanishes at low-$k$: $\omega_{k}^{2}\sim(a\bm{k})^{2}$ [70] we obtain the two limiting cases:
@@
That is, the large length scale temperature is controlled by $T_{com}$ since the energy created on small length scales by the local bath is damped by $\gamma_{com}$ on every scale. While the small length scale temperature is controlled by both the global and local bath. In the following, we will assume that the system is more weakly coupled to the local bath compared to the global bath: $\gamma_{loc}\gg\gamma_{com}$ which leads to $\tilde{T}_{k\to\infty}=T_{loc}$.
The equilibrium-like nature of our equation stems from the assumption that both baths are delta correlated and that each mode is independent on the others in $k$ space. This would not be the case in a realistic system with anharmonic terms for instance and could potentially disrupt the picture given in this section. We discuss the case of an anharmonic lattice in Appendix C and argue that the phenomenology described here should not change qualitatively.
@@ II.2 Hyperuniformity
As already noted in Ref. 33, a key requirement to obtain a long-range ordered 2D crystal is hyperuniformity i.e. the suppression of density fluctuations at arbitrarily large length scales. Within the context of the proposed theoretical model, we derive here the analytical expression of the structure factor $S(\bm{k})$ whose vanishing low-$\bm{k}$ limit indicates the presence of hyperuniformity in a system [34].
@@
When $T_{com}=0$ and $\gamma_{com}\neq 0$, since the SDF is finite at low-$k$, the structure factor behaves as $S(\bm{k})\sim\bm{k}^{2}$, unveiling the hyperuniformity of our non-equilibrium system when the COM position or equivalently the momentum is conserved by the local noise at finite global damping. This aligns with results found in Refs. 73, 50, 74, 75, 76, 33, 77 and confirms that our model includes a suitable limit to study the emergence of long-range order in dimension lower than three.
@@
where $\bm{G}$ is one of the inner Bragg-peak vectors of the crystal and the last equality follows from the gaussianity of the stationary probability distribution function. In a long-range ordered crystal, the correlation function decays to a constant value while in a quasi-long-range 2D equilibrium crystal it exhibits an algebraic decay [4, 7, 79, 80, 81, 82]. From Eq. (14), it is clear that the asymptotic behavior of the displacement correlation function $\langle(\bm{u}_{n}-\bm{u}_{m})^{2}\rangle$ governs the translational long-range behavior of the system. Specifically, if the displacement correlation function reaches a constant or decays at large distances, true long-range order is expected, otherwise, only quasi-long-range order or short-range order is found. In the infinite system size limit and at large distances compared to the lattice spacing it can be approximated as (see Appendix D) :
@@
where the angle average of the dot product appearing in Eq. (14) was taken into account by assuming an isotropic crystal. At non zero $T_{com}$, the large-scale decay of the correlation function is algebraic indicating quasi-long-range order [83]. Nonetheless, in contrast to the equilibrium picture, the average kinetic energy of the system can be arbitrarily high, the only crucial factor for the translational quasi-long-range behavior of the system at large scale, is the temperature of the COM or equivalently, the temperature of the phonons existing on these scales. We also note that, as expected when the COM is conserved at finite global damping ($T_{com}=0$ and $\gamma_{com}\neq 0$), the system displays genuine long-range order as observed in Ref. 33 and the HMW theorem is broken.
@@
with $\beta>0$. Moreover, the relation between the structure factor and long-range order is not limited to the particular mechanism of hyperuniformity we present in this work but is general and applies as well for different hyperuniform systems such as the critical point of absorbing phase transition [75].
@@ II.4 Breakdown of the HMW theorem
It is interesting to note that hyperuniformity of equilibrium crystals is solely achieved through long-range or external interactions [34], in contrast with our system and that of others, where hyperuniformity is allowed through the breakdown of the FDT and correlated noise [24, 84, 73, 85, 86]. This is reminiscent of long-range correlations induced by bulk conservation law and dissipation in self-organized criticality [87, 88, 44, 54, 89, 90, 91, 92, 93, 94]. See Ref. 95 for a recent review. Moreover, COM conservation is not the only route to hyperuniformity and systems with multiplicative [96, 84] or time correlated noise can have suppressed long-range fluctuations [36, 24]. Additionally, recent work on particular non-equilibrium collective excitations in active solids, known as entropons [69, 97], highlight that on a general basis, active forces only add an additional term to the SDF, linked to entropy production but do not affect its phononic part. Thus, active forces alone cannot tame the $1/k^{2}$ divergence of the phononic part of the equilibrium SDF and in general, increase the density fluctuations [30, 28, 29]. Indeed, active short-ranged systems exhibiting hyperuniformity usually lose this property in the presence of a global thermal equilibrium-like noise [98, 27, 99, 37]. This supports the idea that non-equilibrium-like colored or multiplicative noises are crucial for hyperuniformity and translational long-range order in 2D crystals featuring short-ranged interactions. Further discussions about entropons in our system can be found in Appendix E. Note however, that hyperuniformity in passive or active scalar field theories has been found even in the presence of a thermal noise for model B type equations [100, 101, 102, 103].
@@ III.1 A coarse grained model
We explore a practical application of our theoretical framework to investigate the breakdown of the HMW theorem and the dependence of (quasi-)long-range order on the noise in realistic non-equilibrium systems. To conduct this investigation, we use the model introduced in Ref. 38 with an additional global noise. This can be thought as an underdamped continuous-time analogue of a random organization model [104] describing sheared suspensions [105]. In the absence of global noise, these models present an absorbing phase transition, are hyperuniform [34] at the transition point [75, 84] and, if the COM or momentum is locally conserved by the collisions, in the active state [74, 77, 73]. As will be further discussed in the conclusion, this model has been used to describe the experimental realization of vibrated quasi-2d granular systems [106, 38, 107].
We study this system by performing hybrid event-driven/time-stepped molecular dynamics simulations [108] (see Appendix F) of $N$ active hard disks of diameter $\sigma$ and mass $m$ in a 2D square of size $L$ with periodic boundary conditions. We define an arbitrary unit of time $\hat{\tau}$. The (granular) temperature $T$ of the system is defined as its global kinetic energy:
@@
where $0\leq\alpha\leq 1$ is the coefficient of restitution, $\Delta>0$ is a velocity injection term, $\bm{v}_{i}^{\prime}$ the post-collision velocity of particle $i$, $\bm{v}_{i}$ its pre-collision velocity, and $\hat{\bm{\sigma}}_{ij}$ and $\bm{v}_{ij}$ are respectively the unit vector joining particles $i$ and $j$ and the relative velocity between them. In the limit $\Delta=0$, the usual collision rule of dissipative hard disks is recovered. This model is similar to the one introduced in Ref. 73 for liquids.
@@
Hence, to recapitulate the differences and similarities between the harmonic crystal presented in the theory and the granular hard disks of interest here: both models feel an external global Gaussian bath at temperature $T_{com}$ and damping $\gamma_{com}$, however, the active collisions in the granular system are modeled as an effective local bath at temperature $T_{loc}$ and damping $\gamma_{com}\omega_{k}^{2}$ in the theory. From a physical standpoint, the effective $T_{loc}$ in the granular system is technically a function of every parameter of our simulation including, $T_{com}$ and $\gamma_{com}$ because the energy change at the collision depends on the relative velocities of the particles colliding, as understood from Eq. (23), which are affected by the global bath and the packing fraction. This will not change the physical picture given in this article nonetheless, we should keep it in mind. This peculiarity is further discussed in Appendix G.
When $T_{com}=0$, the collision rule given by Eq. (23) locally conserves the momentum and the position of the COM. Moreover, the global damping Eq. (21) damps every modes indistinctly. Hyperuniformity is thus expected [73] and we recover the system studied in Ref. 38 with an absorbing state at low density (a feature absent in the lattice model because $T_{loc}$ is independent on $T_{com}$ and $\gamma_{com}$). Indeed, at low density or packing fraction $\phi=N\pi\sigma^{2}/4L^{2}$, the active collisions cannot compensate for the energy dissipated by the drag during the free flight leading the system to an arrested state. Denser systems attain an active steady state where the energy injection at collision is compensated by the dissipation of the damping.